OpenStudy (anonymous):

What is the height of a box that has a volume of 26.35 cubic centimeters and a base area of 4 square centimeters?

OpenStudy (anonymous):

h * x^2 = 125 (base height width, base=width so x^2) h= 125/x^2 base cost = $5*x^2 side and top cost = $2*(4*(h * x) + $2*(x^2) total = 5x^2 + 2*(4*(h * x) + 2*(x^2) T= 5x^2 + 8hx +2x^2 T= 7x^2+ 8hx T= 7x^2+8x(125/x^2) T= 7x^2 +1000/x